### v1

@@ C.3 Length Correction
Appendix D Examples and Case Study
Tables 7 to 8 illustrate a classic conversation extracted from A Storm of Swords (A Song of Ice and Fire, #3) and the corresponding simulation by CoSER 70B. In CoSER 70B’s simulation, when confronted with Sansa Stark, Lysa Arryn utters her iconic line, “Grown enough to be wed, wed enough to be bedded”, reflecting her personality and worldview. This indicates that CoSER models excellently recall and apply character-related knowledge from their pretrained data.
Tables 9 to 12 present another scene from A Dance with Dragons (A Song of Ice and Fire, #5), specifically the walk of atonement, in which Cersei Lannister is forced to walk naked through the streets, facing both physical and mental humiliation while striving to preserve her dignity. We present the original dialogue alongside simulations by CoSER 70B, GPT-4o, and Claude-3.5-Sonnet. Notably, CoSER 70B faithfully captures the suppressed anger of Cersei Lannister as depicted in the original conversation, whereas the other models, including GPT-4o and Claude-3.5-Sonnet, resort to a stereotypical portrayal of her arrogance and pride.
Simulation Case (from A Storm of Swords (A Song of Ice and Fire, #3))
Simulation Case (from A Storm of Swords (A Song of Ice and Fire, #3))
Simulation Case (from A Dance with Dragons (A Song of Ice and Fire, #5))
Simulation Case (from A Dance with Dragons (A Song of Ice and Fire, #5))
Simulation Case (from A Dance with Dragons (A Song of Ice and Fire, #5))
Simulation Case (from A Dance with Dragons (A Song of Ice and Fire, #5))
Appendix E Additional Results
@@
Table  15 shows experiment results when multi-agent systems continue conversations from the first $k=3$ original messages.
Tables  15 presents detailed evaluation results where our multi-agent simulations start from the first $k=3$ original messages.
@@
Appendix F Prompts
@@
In this section, we list the detailed prompts for: 1) dataset curation in Tables  18 to  20; 2) RPLA and multi-agent simulation in Tables  21 to  22, which have been carefully optimized based on our experience in multi-agent simulation; 3) Penalty-based LLM Judging in Tables  23 to  24.

### v2

@@ C.3 Length Correction
C.4 Human Evaluation
The alignment metrics are measured via the frequency of agreement between LLM and human judges when comparing model pairs, excluding cases where judges assign similar scores to both models (i.e., score differences $\leq$ 1 for human judges or $\leq$ 5% for automatic metrics).
Appendix D Additional Results
Evaluation with Different LLMs as Judges
We conduct additional experiments using alternative judge models for GCA evaluation. We evaluate 7 representative models as shown in Table 11, using three different judge models: GPT-4o, DeepSeek-R1, and DeepSeek-V3. This comparison aims to examine whether our conclusions remain consistent across different LLM judges.
Model GPT-4o DeepSeek-R1 DeepSeek-V3
GPT-3.5 52.8 35.9 40.5
LLaMA-3.1-8B 51.8 37.2 36.8
Abab7-preview 53.7 41.5 40.4
CoSER-8B 56.1 44.5 45.9
GPT-4o 58.5 48.4 46.1
Claude-3.5-Sonnet 56.2 54.8 40.7
CoSER-70B 57.4 50.8 47.7
Table 11: Scores (%) assigned by different LLM judges. Results demonstrate consistent superiority of CoSER models, while revealing judge-specific preferences.
The results in Table 11 reveal several important findings: 1) Despite these variations in absolute scores, the relative ranking patterns remain largely consistent, with CoSER models maintaining competitive or superior performance across all settings. 2) Different judges exhibit distinct preferences, confirming potential bias in LLM-based evaluation. While GPT-4o shows preference for GPT models, DeepSeek-R1 and DeepSeek-V3 favor CoSER-70B and Claude-3.5-Sonnet over GPT-4o.
@@ Appendix D Additional Results
Table  14 shows experiment results when multi-agent systems continue conversations from the first $k=3$ original messages.
Tables  14 presents detailed evaluation results where our multi-agent simulations start from the first $k=3$ original messages.
@@
Comparing GCA Simulation with AI-driven Storytelling Methods
To provide a more comprehensive evaluation against prior work in digital actors and AI-driven storytelling, we compare our GCA simulation approach with HollmWood (Chen et al., 2024c), a representative method for AI-generated storytelling. We conduct this comparison using 30 samples from CoSER Test, employing GPT-4o as both the actor and writer in both frameworks to ensure fair comparison. The generated stories and character interactions are evaluated using GCA evaluation. Results demonstrate that GCA simulation achieves an average score of 59.2%, significantly outperforming HollmWood’s 50.2%. These findings highlight that our given-circumstance acting approach produces more authentic and human-like character interactions compared to previous AI storytelling methods. The superior performance of GCA simulation can be attributed to its focus on character-driven dialogue generation within well-defined circumstances, rather than plot-driven narrative construction, leading to more faithful character portrayals and realistic conversational dynamics.
Appendix E Authentic v.s. LLM-Generated Role-playing Data
To demonstrate the superiority of authentic role-playing data over synthesized alternatives, we conduct two complementary experiments comparing authentic conversations from CoSER dataset with LLM-generated dialogues.
GCA Evaluation of Dialogue Quality
We apply GCA evaluation to score authentic (groundtruth) and LLM-generated dialogues on 100 samples from CoSER Test without providing groundtruth to LLM judges as reference. We use DeepSeek-R1 as the judge model due to its superior alignment with human evaluators (as shown in §5.3). GPT-4o is used to generate synthetic dialogues via GCA simulation. Results show that authentic dialogues from CoSER achieve a score of 85.1%, significantly outperforming GPT-4o generated dialogues at 76.2%.
Training Effectiveness Comparison
We further examine the training effectiveness of authentic versus synthetic dialogues. We fine-tune LLaMA-3.1-8B using authentic or LLM-generated dialogues for 200 samples from CoSER Test for 4 epochs, using the same settings as CoSER-8B. We evaluate the two trained models via GCA evaluation on 100 new samples. The model trained on authentic dialogues from CoSER achieves 55.3% score, while the model trained on GPT-4o generated dialogues achieves 54.7%.
These results demonstrate that authentic dialogues consistently outperform LLM-generated alternatives in both intrinsic quality and training effectiveness. These findings highlight the importance of authentic, high-quality role-playing data for developing effective RPLAs.
Appendix F Examples and Case Study
Tables 17 to 18 illustrate a classic conversation extracted from A Storm of Swords (A Song of Ice and Fire, #3) and the corresponding simulation by CoSER 70B. In CoSER 70B’s simulation, when confronted with Sansa Stark, Lysa Arryn utters her iconic line, “Grown enough to be wed, wed enough to be bedded”, reflecting her personality and worldview. This indicates that CoSER models excellently recall and apply character-related knowledge from their pretrained data.
Tables 19 to 22 present another scene from A Dance with Dragons (A Song of Ice and Fire, #5), specifically the walk of atonement, in which Cersei Lannister is forced to walk naked through the streets, facing both physical and mental humiliation while striving to preserve her dignity. We present the original dialogue alongside simulations by CoSER 70B, GPT-4o, and Claude-3.5-Sonnet. Notably, CoSER 70B faithfully captures the suppressed anger of Cersei Lannister as depicted in the original conversation, whereas the other models, including GPT-4o and Claude-3.5-Sonnet, resort to a stereotypical portrayal of her arrogance and pride.
Table 23 presents an output example of LLM critics, demonstrating the identified performance flaws in GCA simulation.
Simulation Case (from A Storm of Swords (A Song of Ice and Fire, #3))
Simulation Case (from A Storm of Swords (A Song of Ice and Fire, #3))
Simulation Case (from A Dance with Dragons (A Song of Ice and Fire, #5))
Simulation Case (from A Dance with Dragons (A Song of Ice and Fire, #5))
Simulation Case (from A Dance with Dragons (A Song of Ice and Fire, #5))
Simulation Case (from A Dance with Dragons (A Song of Ice and Fire, #5))
Examples of LLM Critiques (from Pride and Prejudice)
Actor Model Claude-3.5-Sonnet
Critic Model GPT-4o
Overall Scores Storyline Consistency: 60.0%  Anthropomorphism: 45.0% Character Fidelity: 80.0%  Storyline Quality: 80.0%
Storyline Consistency Flaws Storyline Consistency (Severity 3): Mr. Bennet’s tone and behavior are less sarcastic and more emotionally engaged than in the original conversation, indicating a change from his characteristic detachment and sarcasm to a more invested and sincere demeanor. Storyline Consistency (Severity 3): Elizabeth expresses her transformation in feelings toward Mr. Darcy and discusses her changed sentiments openly with Mr. Bennet. This level of openness and emotional vulnerability is a departure from the original scenario. Storyline Consistency (Severity 4): The simulated conversation includes an emotional resolution with Elizabeth expressing affection and admiration for Mr. Darcy, which contrasts with the more restrained and ambiguous end of the original conversation. Storyline Consistency (Severity 4): Mr. Bennet shows a deep level of affectionate concern and engages in a sentimental exchange with Elizabeth, inconsistent with the original conversation where his reaction focuses more on amusement and irony.
Anthropomor- phism Flaws Emotional Depth (Severity 3): Elizabeth speaks aloud all her thoughts and emotions, lacking the subtlety she usually employs. Persona Coherence (Severity 4): Mr. Bennet’s reactions and speech lack the biting sarcasm characteristic of his persona, showing more emotional openness and lacking the irony that defines him. Self-identity (Severity 3): Elizabeth’s willingness to openly share her entire emotional journey regarding Mr. Darcy doesn’t align with her usually more guarded nature. Emotional Depth (Severity 4): Elizabeth’s extensive exposition of her transformation of feelings lacks the psychological complexity and indirect communication she often displays.
Character Fidelity Flaws Personality & Behavior (Severity 4): Elizabeth openly admits her feelings about Mr. Darcy to her father, discussing her transformation of feelings and admiration for him. Character Language (Severity 3): Mr. Bennet shows deep concern about Elizabeth’s feelings and speaks with a tenderness not typical of his character in the book. Relationship & Social Status (Severity 3): The interaction between Mr. Bennet and Elizabeth is unusually emotional and tender compared to their typically sardonic and witty exchanges.
Storyline Quality Flaws Logical Consistency (Severity 3): Elizabeth openly admits her feelings about Mr. Darcy without the preceding tension and subtlety that would naturally lead to such a revelation. The original scenario builds up to her realization with more internal reflection rather than direct admission. Flow & Progression (Severity 3): The emotional exchange and direct discussion about Elizabeth’s feelings for Mr. Darcy happen quickly without sufficient development or anticipation. This leap in intimacy and vulnerability, coupled with Mr. Bennet’s uncharacteristically supportive demeanor, disrupts the story’s traditional emotional pacing.
Table 23: Example of GCA evaluation critique by GPT-4o judge on Claude-3.5-Sonnet’s simulation of a scenario from Pride and Prejudice. The critique demonstrates the detailed flaw identification across multiple evaluation dimensions.
Appendix G Prompts
@@
In this section, we list the detailed prompts for: 1) dataset curation in Tables  24 to  26; 2) RPLA and multi-agent simulation in Tables  27 to  28, which have been carefully optimized based on our experience in multi-agent simulation; 3) Penalty-based LLM Judging in Tables  29 to  30.